Royal Queen Petroleum has the following data for its Bayou Field: Property cost (acquisition cost)....... $ 60,000 Drilling cost (one well)................... 280,000 Estimated selling cost per bbl..................80 Estimated lifting cost per bbl...................26 State severance tax..................................5% Royalty interest...................................12.5% The company is considering two drilling plans which are estimated to have the following production: Well A: 600 bbl per month. Completion cost, $500,000. Well B: 1,000 bbl per month. Completion cost, 800,000.

Required: a. Determine the number of months needed for payout on each plan.

b. If the company depends on the payback method for its investment decision, which plan will be more preferred?
Answer #1
Please give positive ratings so I can keep answering. If you have any queries please comment. Thanks!
Royal Queen Petroleum Amount $
Particulars Well A Well B
Property cost      60,000.00         60,000.00
Drilling cost 280,000.00       280,000.00
Completion cost 500,000.00       800,000.00
Total cash outflow 840,000.00 1,140,000.00 A
Selling cost per bbl             80.00                 80.00 B
Estimated lifting cost per bbl            (26.00)                (26.00) C
State severance tax @ 5%              (4.00)                  (4.00) D
Royalty interest @ 12.5%            (10.00)                (10.00) E
Net income per bbl             40.00                 40.00 F=B+C+D+E
Estimated production per month           600.00            1,000.00 G
Net income per month     24,000.00         40,000.00 H=F*G
Payback Period (months)             35.00                 28.50 I=A/H
Plan B will be more preferred as its payback period is less.